# Imputation of Missing Values in Multi-View Data

### v3

@@ 2 Methods
To prevent wasting observed data, missing values can be imputed. The simplest form of imputation is to simply replace each missing value with a constant. A very common choice is the unconditional mean of the feature, a procedure known as (unconditional) mean imputation (MI). If one is primarily interested in prediction, MI has some favorable properties: Its computational cost is extremely small, and it has been shown that MI is universally consistent for prediction even for MAR data, as long as the learning algorithm used is also universally consistent [19]. Here consistent means that, given an infinite amount of training data, the prediction function achieves the error rate of the best possible prediction function (i.e. the Bayes rate), while universal means that the procedure is consistent for all possible data distributions [19]. However, MI is often criticised because it is known to distort the data distribution by attenuating existing correlations between the features, underestimating the variance, and causing bias in almost any estimate other than the mean [18].
@@
We can also categorize the existing imputation methods depending on whether they perform unconditional or conditional imputation. We define an unconditional imputation method as any method in which the imputation of a missing value is based solely on other observations of the same feature, that is, the imputation takes place within a single column of the feature matrix. The aforementioned mean imputation is a classical example of an unconditional imputation method. By contrast, a conditional imputation method is any method in which the imputation of a missing value is based, in part or completely, on observations of other features, that is, the imputation uses different columns of the feature matrix. Most sophisticated imputation methods, such as Bayesian multiple imputation and PMM, are conditional imputation methods. The distinction between unconditional and conditional imputation methods is of particular interest for feature selection. Unconditional imputation methods, such as mean substitution, use only the univariate distributions for imputation, so that the imputed feature remains in some sense ‘pure’ and free from contamination from other features. However, as mentioned earlier, mean substitution is known to distort the data distribution by attenuating existing correlations between the features [18].
By contrast, some (but not all) conditional imputation methods preserve the correlations between features [18]. However, in this case the imputed values depend on other features in the data. In the event that a selected feature has a large number of imputed values, this may lead to difficulties in interpretation, since a large proportion of the selected feature is derived from other features. Nevertheless, a recent study on the effect of imputation methods on feature selection suggests sophisticated conditional imputation methods generally lead to better results than unconditional imputation [33]. Because it is not possible to both perform the imputation independent of other features and preserve existing correlations, one has to choose between one or the other.
@@ 2.2 Multi-View Stacking and the StaPLR Algorithm
StaPLR denotes the special case of MVS where all learners are penalized logistic regression learners [14]. In particular, we choose $A_{\text{b}}$ to be a $L_{2}$-penalized logistic regression learner including cross-validation for the tuning parameter, so that each $\hat{f}_{v}$ is a fully tuned logistic ridge regression classifier [14]. For $A_{\text{m}}$ we choose a $L_{1}$-penalized nonnegative logistic regression learner including cross-validation for the tuning parameter, so that $\hat{f}_{\text{meta}}$ is a fully tuned logistic nonnegative lasso classifier [14, 15]. In this setup, the base-learners induce shrinkage within each view, while the meta-learner selects or discards views based on their predictive relevance [14]. StaPLR, and the MVS model in general, share some similarities with the recently proposed Mixture-of-Views (MoV) model [69]. Both are based on the “divide and conquer” paradigm of dividing a complex problem into simpler ones and then combining them [69]. However, the MoV model is a multi-view generalization of the mixture-of-experts model (MoE) [70, 71, 72], while StaPLR/MVS is a multi-view generalization of stacking [73]. This means there are two important differences between the models. The first is that in the MoV model, the weights are assigned to each of the view-specific classifiers based on within-sample performance, while in StaPLR/MVS they are assigned based on estimated out-of-sample performance. Basing the weights on estimated out-of-sample performance is the core idea behind stacking, and can filter out bias and improve generalization performance [73]. The second difference is that in the MoV model the goal is to find the most predictive set of views for each individual sample, while in StaPLR the goal is to find the most predictive set of views across all samples. In our case, only the second type of view selection is of interest as it has the potential to reduce data collection in future experiments.
@@
2       $\hat{f}_{v}=A_{\text{b}}(\bm{X}^{(v)},\bm{y})$ foreach $k$ = 1 to $K$ do
3             $\hat{f}_{v,k}=A_{\text{b}}(\bm{X}^{(v)}_{i\notin S_{k}},\bm{y}_{i\notin S_{k}})$ $\bm{z}^{(v)}_{k}=\hat{f}_{v,k}(\bm{X}^{(v)}_{i\in S_{k}})$
@@ 2       $\hat{f}_{v}=A_{\text{b}}(\bm{X}^{(v)},\bm{y})$ foreach $k$ = 1 to $K$ do
$\bm{Z}=(\bm{z}^{(1)},\dots,\bm{z}^{(V)})$ $\hat{f}_{\text{meta}}=A_{\text{m}}(\bm{Z},\bm{y})$ $\hat{f}_{\text{stacked}}=\hat{f}_{\text{meta}}\circ(\hat{f}_{\text{1}}\dots\hat{f}_{\text{V}})$
@@ 2.3 The Proposed Imputation Method
As shown in Algorithm 1 and Figure 1, StaPLR performs view selection through two steps: a dimension reduction step, and a feature selection step. First, a matrix of cross-validated predictions $\bm{Z}$ is constructed, in which each view is represented by a single column. The meta-learning algorithm then performs feature selection in this reduced space, which translates to view selection in the original feature space. Note that the reduction of dimensionality is generally very large: If the complete concatenated feature matrix $\bm{X}=(\bm{X}^{(1)},\dots,\bm{X}^{(V)})$ has dimensionality $n\times m$, then $\bm{Z}$ has dimensionality $n\times V$, and generally $V\ll m$. This further implies that, while the original data is usually very high dimensional, this is often not the case in the reduced space; see for example van Loon et al. [16] for an application of StaPLR where $n\ll m$, but $n>V$. Obviously, it is much less computationally intensive to fit models in the reduced space, than it is to fit them in the original space. We therefore propose to impute any missing values in this reduced space rather than in the original feature space. Because of the difference in dimensionality, this is computationally very attractive. Furthermore, as long as we assume $n>V$, imputation methods which would perform poorly in the high-dimensional feature space can instead be applied in the reduced space.
@@ 3 Empirical Evaluation
The first imputation algorithm is simple unconditional mean imputation (MI), where each missing value is replaced by the mean of the observed values of the feature. The second algorithm is predictive mean matching (PMM) [18], a state-of-the-art multiple imputation algorithm implemented as the default in the R package mice [23]. PMM is based on Bayesian imputation under the normal linear model [18, 74, 75], but instead of using a Bayesian draw of the regression coefficient to directly calculate an imputed value, it is used to find a set of $d$ candidate ‘donors’, which are the $d$ observations for which the predicted value under the imputation model is closest to the value drawn for the missing observation [18, 23]. The imputed value is then a random draw from the observed values of the donors. Advantages of PMM over direct Bayesian imputation include that the method allows for discrete target variables, the imputed values can never go outside the range of the observed data, and the method is more robust to model misspecification [18]. Disadvantages compared to other imputation methods include the fact that PMM tends to work poorly in high-dimensional settings, and that the required computation time is so large that we can only apply PMM at the meta-level, as feature-level imputation using PMM proved entirely infeasible (see A). When applying PMM, we use the default value of 5 for both the number of candidate donors and the number of imputations. With multiple imputation comes the need to somehow combine the results of the different imputed data sets. We evaluate two strategies for combining the different imputed data sets: The first strategy is to generate the matrix of cross-validated predictions $\bm{Z}$ once, use the imputation model to generate 5 imputations, and then take the average of the imputed values to train the meta-learner. The second strategy is to generate the matrix $\bm{Z}$ 5 times, each time using different cross-validation folds, and then impute each matrix once before averaging the matrices. The benefit of the second strategy is that it also averages over different ways in which observations are allocated to cross-validation folds, but it is approximately 5 times more computationally expensive. To distinguish between the two strategies, we will refer to them as mPMM and cvPMM, respectively.
@@ 4.1 View Selection
In terms of the TPR (Table 2), CCA performs worse than all imputation methods. Both MI and mMI perform well in terms of TPR when compared with the other imputation methods. The mMF, mPMM and cvPMM method again show similar performance. The observed differences between MF and mMF do not appear to follow a consistent pattern.
In terms of the FPR (Table 3), a pattern can be observed where mean imputation (MI and mMI) performs really well when the view to be imputed corresponds to noise, but performs poorly when the view to be imputed corresponds to signal. The performance of MF shows the opposite behavior: it performs better when the view to be imputed corresponds to signal than when it corresponds to noise. Similar behavior can be observed for its meta-level implementation mMF. MOFA imputation also shows this pattern but in an extreme fashion: When the missing view correponds to noise the FPR rises dramatically, up to 0.90. Comparing MF and mMF shows that MF has a slightly lower FPR when the size of the missing view is 5, but has a considerably higher FPR than mMF when the size of the missing view is 5000. The observed differences in FPR between mMF and mPMM and cvPMM are small under 50% missingness. Under 90% missingness larger differences are observed when the signal view with 5 features is missing, and when the noise view with 5000 features is missing. In terms of FPR, CCA performs well when compared to the imputation methods.
The relative performance of the imputation methods in terms of FDR (Table 4) is similar to that in terms of FPR. CCA also performs fairly well in terms of FDR, just like it does in terms of FPR.
@@
Table 1: Mean proportion of correctly selected views. CDA = complete data analysis; CCA = complete case analysis; MI = feature-level mean imputation; MF = feature-level missForest; mMI = meta-level mean imputation; mMF = meta-level missForest; mPMM = meta-level predictive mean matching, generating $\bm{Z}$ once; cvPMM = meta-level predictive mean matching, generating $\bm{Z}$ five times; MOFA = multi-factor omics analysis imputation.
@@
Table 2: True positive rate (TPR). CDA = complete data analysis; CCA = complete case analysis; MI = feature-level mean imputation; MF = feature-level missForest; mMI = meta-level mean imputation; mMF = meta-level missForest; mPMM = meta-level predictive mean matching, generating $\bm{Z}$ once; cvPMM = meta-level predictive mean matching, generating $\bm{Z}$ five times; MOFA = multi-factor omics analysis imputation.
@@
Table 3: False positive rate (FPR). CDA = complete data analysis; CCA = complete case analysis; MI = feature-level mean imputation; MF = feature-level missForest; mMI = meta-level mean imputation; mMF = meta-level missForest; mPMM = meta-level predictive mean matching, generating $\bm{Z}$ once; cvPMM = meta-level predictive mean matching, generating $\bm{Z}$ five times; MOFA = multi-factor omics analysis imputation.
@@
Table 4: False discovery rate (FDR). CDA = complete data analysis; CCA = complete case analysis; MI = feature-level mean imputation; MF = feature-level missForest; mMI = meta-level mean imputation; mMF = meta-level missForest; mPMM = meta-level predictive mean matching, generating $\bm{Z}$ once; cvPMM = meta-level predictive mean matching, generating $\bm{Z}$ five times; MOFA = multi-factor omics analysis imputation.
@@ 4.2 Computation Time
Figure 7 shows the log computation time for all imputation methods, as well as CDA and CCA, with 90% missingness. Figure 8 shows the log computation time with 50% missingness. Note that the computation times reported here capture both imputation and model fitting. It can be observed that CCA has the lowest computation time across all plots, since it constitutes a reduction in the amount of observations in the training data. By comparing MI with mMI, and MF with mMF, it can be observed that meta-level imputation is faster than base-level imputation. This is particularly noticeable when the view to be imputed is large (i.e. when $V_{4}$ is missing), when MF is the slowest method by a large margin. cvPMM is considerably slower than mPMM, and is the slowest method when the view to be imputed is small (i.e. when $V_{1}$ is missing).
@@ 4.3 Summary of Simulation Results
  • Meta-level imputation performs approximately as well as feature-level imputation in terms of Accuracy, TPR, FPR, or FDR.
@@ 5 Application to multi-view MRI data
It can be observed in Figures 9 and 10 that mean imputation, either at the base level or the meta level, performs very well if the functional scan is missing, even performing slightly better than the complete data. However, it performs poorly when the structural scan is missing. Meta-level missForest imputation peforms better than meta-level PMM, especially when the functional scan is missing. In real data the ‘true’ underlying model is unknown and quantities such as the TPR, FPR and FDR cannot be calculated. However, we can compare the coefficients obtained from fitting the model on the imputed data with those obtained from fitting the model on the complete data (Figure 11). Again it can be observed that mean imputation performs poorly when the structural scan is missing, as it attenuates the coefficients of the most important views (grey matter density and cortical thickness) while inflating those of comparatively less important views (such as radial diffusivity and ALFF). This behavior is comparable to that observed in our simulation study (Section 4.1), although in the real data it does lead to a substantial reduction in accuracy (Figure 9). Using meta-level PMM when the functional scan is missing leads to many additional functional views being selected, which is also associated with a reduction in accuracy (Figure 9). Out of all the imputation methods, meta-level missForest imputation leads to coefficients which are on average closest to those obtained from the complete data (Figure 11). In addition, it can be observed in Figure 12 that it is also faster than meta-level PMM.

### v4

# Stacked Penalized Logistic Regression for Selecting Views in Multi-View Learning

@@ Abstract
keywords missing data, imputation, multi-view learning, stacked generalization, feature selection
@@ 2 Methods
To prevent wasting observed data, missing values can be imputed. The simplest form of imputation is to replace each missing value with a constant. A very common choice is the unconditional mean of the feature, a procedure known as (unconditional) mean imputation (MI). If one is primarily interested in prediction, MI has some favorable properties: Its computational cost is extremely small, and it has been shown that MI is universally consistent for prediction even for MAR data, as long as the learning algorithm used is also universally consistent [19]. Here consistent means that, given an infinite amount of training data, the prediction function achieves the error rate of the best possible prediction function (i.e., the Bayes rate), while universal means that the procedure is consistent for all possible data distributions [19]. However, MI is often criticized because it is known to distort the data distribution by attenuating existing correlations between the features, underestimating the variance, and causing bias in almost any estimate other than the mean [18].
@@
We can also categorize the existing imputation methods depending on whether they perform unconditional or conditional imputation. We define an unconditional imputation method as any method in which the imputation of a missing value is based solely on other observations of the same feature, that is, the imputation takes place within a single column of the feature matrix. The aforementioned mean imputation is a classic example of an unconditional imputation method. By contrast, a conditional imputation method is any method in which the imputation of a missing value is based, in part or completely, on observations of other features, that is, the imputation uses different columns of the feature matrix. Most sophisticated imputation methods, such as Bayesian multiple imputation and PMM, are conditional imputation methods. The distinction between unconditional and conditional imputation methods is of particular interest for feature selection. Unconditional imputation methods, such as mean imputation, use only the univariate distributions for imputation, so that the imputed feature remains in some sense ‘pure’ and free from contamination from other features. However, as mentioned earlier, mean imputation is known to distort the data distribution by attenuating existing correlations between the features [18].
By contrast, some (but not all) conditional imputation methods preserve the correlations between features [18]. However, in this case the imputed values depend on other features in the data. In the event that a selected feature has a large number of imputed values, this may lead to difficulties in interpretation, since a large proportion of the selected feature is derived from other features. Nevertheless, a recent study on the effect of imputation methods on feature selection suggests sophisticated conditional imputation methods generally lead to better results than unconditional imputation methods [33]. Because it is not possible to both perform the imputation independent of other features and preserve existing correlations, one has to choose between one or the other.
@@ 2.2 Multi-View Stacking and the StaPLR Algorithm
StaPLR denotes the special case of MVS where all learners are penalized logistic regression learners [14]. In particular, we choose $A_{\text{b}}$ to be an $L_{2}$-penalized logistic regression learner including cross-validation for the tuning parameter, so that each $\hat{f}_{v}$ is a fully tuned logistic ridge regression classifier [14]. For $A_{\text{m}}$ we choose an $L_{1}$-penalized nonnegative logistic regression learner including cross-validation for the tuning parameter, so that $\hat{f}_{\text{meta}}$ is a fully tuned logistic nonnegative lasso classifier [14, 15]. In this setup, the base-learners induce shrinkage within each view, while the meta-learner selects or discards views based on their predictive relevance [14]. StaPLR, and the MVS model in general, share some similarities with the recently proposed Mixture-of-Views (MoV) model [69]. Both are based on the “divide and conquer” paradigm of dividing a complex problem into simpler ones and then combining them [69]. However, the MoV model is a multi-view generalization of the mixture-of-experts model (MoE) [70, 71, 72], while StaPLR/MVS is a multi-view generalization of stacking [73]. This means there are two important differences between the models. The first is that in the MoV model, the weights are assigned to each of the view-specific classifiers based on within-sample performance, while in StaPLR/MVS they are assigned based on estimated out-of-sample performance. Basing the weights on estimated out-of-sample performance is the core idea behind stacking, and can filter out bias and improve generalization performance [73]. The second difference is that in the MoV model the goal is to find the most predictive set of views for each individual sample, while in StaPLR the goal is to find the most predictive set of views across all samples. In our case, only the second type of view selection is of interest as it has the potential to reduce data collection in future experiments.
@@
2       $\hat{f}_{v}=A_{\text{b}}(\bm{X}^{(v)},\bm{y})$
3       foreach $k$ = 1 to $K$ do
4             $\hat{f}_{v,k}=A_{\text{b}}(\bm{X}^{(v)}_{i\notin S_{k}},\bm{y}_{i\notin S_{k}})$
5             $\bm{z}^{(v)}_{k}=\hat{f}_{v,k}(\bm{X}^{(v)}_{i\in S_{k}})$
@@
[…]
10 $\hat{f}_{\text{meta}}=A_{\text{m}}(\bm{Z},\bm{y})$
$\hat{f}_{\text{stacked}}=\hat{f}_{\text{meta}}\circ(\hat{f}_{\text{1}}\dots\hat{f}_{\text{V}})$
@@ 2.3 The Proposed Imputation Method
As shown in Algorithm 1 and Figure 1, StaPLR performs view selection through two steps: a dimension reduction step, and a feature selection step. First, a matrix of cross-validated predictions $\bm{Z}$ is constructed, in which each view is represented by a single column. The meta-learning algorithm then performs feature selection in this reduced space, which translates to view selection in the original feature space. Note that the reduction of dimensionality is generally very large: If the complete concatenated feature matrix $\bm{X}=(\bm{X}^{(1)},\dots,\bm{X}^{(V)})$ has dimensionality $n\times m$, then $\bm{Z}$ has dimensionality $n\times V$, and generally $V\ll m$. This further implies that, while the original data is usually very high-dimensional, this is often not the case in the reduced space; see for example van Loon et al. [16] for an application of StaPLR where $n\ll m$, but $n>V$. Obviously, it is much less computationally intensive to fit models in the reduced space, than it is to fit them in the original space. We therefore propose to impute any missing values in this reduced space rather than in the original feature space. Because of the difference in dimensionality, this is computationally very attractive. Furthermore, as long as we assume $n>V$, imputation methods which would perform poorly in the high-dimensional feature space can instead be applied in the reduced space.
@@ 3 Empirical Evaluation
The first imputation algorithm is unconditional mean imputation (MI), where each missing value is replaced by the mean of the observed values of the feature. The second algorithm is predictive mean matching (PMM) [18], a state-of-the-art multiple imputation algorithm implemented as the default in the R package mice [23]. PMM is based on Bayesian imputation under the normal linear model [18, 74, 75], but instead of using a Bayesian draw of the regression coefficient to directly calculate an imputed value, it is used to find a set of $d$ candidate ‘donors’, which are the $d$ observations for which the predicted value under the imputation model is closest to the value drawn for the missing observation [18, 23]. The imputed value is then a random draw from the observed values of the donors. Advantages of PMM over direct Bayesian imputation include that the method allows for discrete target variables, the imputed values can never go outside the range of the observed data, and the method is more robust to model misspecification [18]. Disadvantages compared to other imputation methods include the fact that PMM tends to work poorly in high-dimensional settings, and that the required computation time is so large that we can only apply PMM at the meta-level, as feature-level imputation using PMM proved entirely infeasible (see A). When applying PMM, we use the default value of 5 for both the number of candidate donors and the number of imputations. With multiple imputation comes the need to somehow combine the results of the different imputed data sets. We evaluate two strategies for combining the different imputed data sets: The first strategy is to generate the matrix of cross-validated predictions $\bm{Z}$ once, use the imputation model to generate 5 imputations, and then take the average of the imputed values to train the meta-learner. The second strategy is to generate the matrix $\bm{Z}$ 5 times, each time using different cross-validation folds, and then impute each matrix once before averaging the matrices. The benefit of the second strategy is that it also averages over different ways in which observations are allocated to cross-validation folds, but it is approximately 5 times more computationally expensive. To distinguish between the two strategies, we will refer to them as mPMM and cvPMM, respectively.
@@ 4.1 View Selection
In terms of the true positive rate (Table 2), CCA performs worse than all imputation methods. Both MI and mMI perform well in terms of TPR when compared with the other imputation methods. The mMF, mPMM and cvPMM method again show similar performance. The observed differences between MF and mMF do not appear to follow a consistent pattern.
In terms of the false positive rate (Table 3), a pattern can be observed where mean imputation (MI and mMI) performs really well when the view to be imputed corresponds to noise, but performs poorly when the view to be imputed corresponds to signal. The performance of MF shows the opposite behavior: It performs better when the view to be imputed corresponds to signal than when it corresponds to noise. Similar behavior can be observed for its meta-level implementation mMF. MOFA imputation also shows this pattern but in an extreme fashion: When the missing view corresponds to noise the FPR rises dramatically, up to 0.90. Comparing MF and mMF shows that MF has a slightly lower FPR when the size of the missing view is 5, but has a considerably higher FPR than mMF when the size of the missing view is 5000. The observed differences in FPR between mMF and mPMM and cvPMM are small under 50% missingness. Under 90% missingness larger differences are observed when the signal view with 5 features is missing, and when the noise view with 5000 features is missing. In terms of FPR, CCA performs well when compared to the imputation methods.
The relative performance of the imputation methods in terms of false discovery rate (Table 4) is similar to that in terms of FPR. CCA also performs fairly well in terms of FDR, just like it does in terms of FPR.
@@
Table 1: Mean proportion of correctly selected views. CDA = complete data analysis; CCA = complete case analysis; MI = feature-level mean imputation; MF = feature-level missForest; mMI = meta-level mean imputation; mMF = meta-level missForest; mPMM = meta-level predictive mean matching, generating $\bm{Z}$ once; cvPMM = meta-level predictive mean matching, generating $\bm{Z}$ five times; MOFA = multi-factor omics analysis imputation. The ‘best’ (highest) values in each row are printed in bold (excluding the values for CDA and CCA since these are not imputation methods).
@@
Table 2: True positive rate (TPR). CDA = complete data analysis; CCA = complete case analysis; MI = feature-level mean imputation; MF = feature-level missForest; mMI = meta-level mean imputation; mMF = meta-level missForest; mPMM = meta-level predictive mean matching, generating $\bm{Z}$ once; cvPMM = meta-level predictive mean matching, generating $\bm{Z}$ five times; MOFA = multi-factor omics analysis imputation. The ‘best’ (highest) values in each row are printed in bold (excluding the values for CDA and CCA since these are not imputation methods).
@@
Table 3: False positive rate (FPR). CDA = complete data analysis; CCA = complete case analysis; MI = feature-level mean imputation; MF = feature-level missForest; mMI = meta-level mean imputation; mMF = meta-level missForest; mPMM = meta-level predictive mean matching, generating $\bm{Z}$ once; cvPMM = meta-level predictive mean matching, generating $\bm{Z}$ five times; MOFA = multi-factor omics analysis imputation. The ‘best’ (lowest) values in each row are printed in bold (excluding the values for CDA and CCA since these are not imputation methods).
@@
Table 4: False discovery rate (FDR). CDA = complete data analysis; CCA = complete case analysis; MI = feature-level mean imputation; MF = feature-level missForest; mMI = meta-level mean imputation; mMF = meta-level missForest; mPMM = meta-level predictive mean matching, generating $\bm{Z}$ once; cvPMM = meta-level predictive mean matching, generating $\bm{Z}$ five times; MOFA = multi-factor omics analysis imputation. The ‘best’ (lowest) values in each row are printed in bold (excluding the values for CDA and CCA since these are not imputation methods).
@@ 4.2 Computation Time
Figure 7 shows the log computation time for all imputation methods, as well as CDA and CCA, with 90% missingness. Figure 8 shows the log computation time with 50% missingness. Note that the computation times reported here capture both imputation and model fitting. It can be observed that CCA has the lowest computation time across all plots, since it constitutes a reduction in the amount of observations in the training data. By comparing MI with mMI, and MF with mMF, it can be observed that meta-level imputation is faster than base-level imputation. This is particularly noticeable when the view to be imputed is large (i.e., when $V_{4}$ is missing), when MF is the slowest method by a large margin. cvPMM is considerably slower than mPMM, and is the slowest method when the view to be imputed is small (i.e., when $V_{1}$ is missing).
@@ 4.3 Summary of Simulation Results
  • Meta-level imputation performs approximately as well as feature-level imputation in terms of accuracy, TPR, FPR, and FDR.
@@ 5 Application to multi-view MRI data
It can be observed in Figures 9 and 10 that mean imputation, either at the base level or the meta level, performs very well if the functional scan is missing, even performing slightly better than the complete data. However, it performs poorly when the structural scan is missing. Meta-level missForest imputation performs better than meta-level PMM, especially when the functional scan is missing. In real data the ‘true’ underlying model is unknown and quantities such as the TPR, FPR and FDR cannot be calculated. However, we can compare the coefficients obtained from fitting the model on the imputed data with those obtained from fitting the model on the complete data (Figure 11). Again it can be observed that mean imputation performs poorly when the structural scan is missing, as it attenuates the coefficients of the most important views (grey matter density and cortical thickness) while inflating those of comparatively less important views (such as radial diffusivity and ALFF). This behavior is comparable to that observed in our simulation study (Section 4.1), although in the real data it does result in a substantial reduction in accuracy (Figure 9). Using meta-level PMM when the functional scan is missing leads to many additional functional views being selected, which is also associated with a reduction in accuracy (Figure 9). Out of all the imputation methods, meta-level missForest imputation leads to coefficients which are on average closest to those obtained from the complete data (Figure 11). In addition, it can be observed in Figure 12 that it is also faster than meta-level PMM.